UT Feb 09 2021




Night Partition

D/E Time - twilight until 20:30 - handover to PEPSI (local)

PEPSI 20:30hr - 03:30hr (local)

D/E Time - 03:30 to twilght (local)

Script Location: /home/lbto/Service_Observing/D_E_Feb_Mar_2021/

Night Plan

  • Configure for Mixed-Mode MODS-1 SX - LBC-Red DX
  • MODS1+LBCR mixed-mode on-sky test (OPK)
    • 0.5-1 hr
    • Run a long MODS observation with MOS-type alignment (i.e., dtheta, dx, dy) and a simultaneous LBCR dithered sequence.
      • Make sure we can stop LBCR, refocus/recollimate, and then load & play a new LBCR OB.
    • priority=high (for an LBTB program), seeing=poor, clouds=don't care
  • Reconfigure for Bino-MODS
  • MODS1 Active Optics parameter change (JMH) - 1 hr
    • MODS1 collimating with ~3 different guide stars (don't care what MODS is looking at)
    • priority=medium-low, seeing=average, clouds=don't care
  • UGC 4079 - MODS Sci Program - filler to reach PEPSI handover time (BR) - 1hr-ish
    • This is filler in case actitivites completed before 20:30 - not high priority
    • Seeing up to 1.5", some clouds OK (even 1-2 mags extinction OK)
    • Script and finder Location: /home/lbto/Service_Observing/D_E_Feb_Mar_2021/Rothberg_Science/Local_Mergers/scripts
    • Target is the nucleus of the galaxy - finder charts included - but target is obvious - change stretch to make sure the center of the galaxy is centered in the slit - these are single nucleus targets - nucleus will be brightest of any sources
    • SpecStd: G191B2b
  • Configure for PEPSI until 03:30
  • Reconfigure at 03:30 (local) to LBC-Bino
  • LBC Pointing Hysteresis and Settling Test (JMH/OPK)
    • needs 1 hour max in any conditions where we are open
    • priority=medium-low (interesting for TMS)
  • LBC TMS active control tests
  • or LBC Science - Solar System target recovery (Veillet)
  • If can't do LBC tests for some reason
    • Configure for MODS Science - IMBH Candidate (Rothberg) - 1.5 hr
      • Seeing 1.2" or better, thin clouds or better
      • Script Location: /home/lbto/Service_Observing/D_E_Feb_Mar_2021/Rothberg_Science/IMBH
    • Poor weather (and can't confugure for LBC LMS testing)
      • MODS-bino Backup - Mergers Dual Grating (Rothberg) - up to 1.5" seeing, up to 1-2 mags clouds
        • Script and finder Location: /home/lbto/Service_Observing/D_E_Feb_Mar_2021/Rothberg_Science/Local_Mergers/scripts
        • NGC 3310 (Visbile 03:30-twilight) - 60 min shutter time
        • NGC 6598 (Visible 03:30-twilight) - 60 min shutter time
        • Targets are the nucleus of the galaxy - finder charts included - but target is obvious - change stretch to make sure the center of the galaxy is centered in the slit - these are single nucleus targets - nucleus will be brightest of any sources
        • Spec Std - Feige 67

Night Log

All times in UT unless otherwise noted.

01:15 Opening to cirrusy skies

Mixed-Mode MODS1+LBCR test

Following notes here: Mixed-mode LBC+LUCI/MODS

Summary - Yes, it works and all the offsets, even the first offset of the "repeat" OB (the LBC OB with dec=-90 that is to be played after the focus/collimation) was made, even though the PCSGUI did not show it.

01:18 David authorizing NONE+LBCR. I ran up both LBCB+LBCR and connected only LBCR. I am running a monocular LBCR-only script and not unclicking anything in the User Interface. Since only LBCR is authorized, it should not be necessary to unclick Blue Channel and Blue Mirror.

01:24 dohybrid, /redonly, /x2 (may be a bit bright, but we'll see). We saw, the 32-sec exposure is saturated. No pupils found! 12-deg twilight 18:54MST

01:58 Trying again. Lots of comatic, fuzzy pupils. Glass, Ambient = 0.98, 1.8 C and outside 0.3 C. The dome has a lot of hot air. DIMM=1.1" (Zen-corrected from AM=1.35) and estimated FPIA seeing is ~3". Shift-S to stop so we can move on.

02:07 LBTtools.camera="REDLBC", load LBC package.

02:08 Play pointing OB. LBCR 020935.
  • ptautoadjust. Sent corrections to IE, CA. dIE, dCA = -4.6, -22.9

02:15 David is authorizing MODS1 and LBCR.
  • Pointing check, ok, but collimation is not. The collimation "star" is a galaxy! Going to another collimation star.
  • Warning: GCS guides, PCS not accepting corrections (something that often shows up when one presses "Stop guiding").
  • Next collimation star is much better.

02:31 We need to tweak up LBCR collimation, since it has been a while. Running dofpia, /redonly here using the preset sent for the MODS collimation. Alternatively, one could slew back to the LBCR pointing field, but the OSA would have to release the syncpreset on the MODS side.
  • dofpia, /redonly, /x2

02:41 Sending sync presets:
  • acqMODS --mods1 --bino MODS_acq.txt
    • sending a detxy rel offset -0.602" 1.267" -0.2 deg, absorbed it.
    • tweaked the centering
  • execMODS --mods1 --bino MODS_obs.txt
  • LBC - playing LBC_initial_5ptcustom_pa90.xml
    • Stopping it after 3 dithers to run dofpia
      • 02:57 dofpia, /redonly
    • Playing the OB, LBC_repeat_5ptcustom_pa90.xml
      • The first offset is not done (should be 25, 0 but PCSGUI says it is -20, 20)
        • The first offset is lumped into the preset, as I recall, so that may explain why.
      • The second offset is correctly given (PCSGUI says 30, 25).
      • Saw the 3rd being sent. Stopped. * Stopped MODS observation.

FILENAME                  OBJECT DITHOFFX DITHOFFY TELRA TELDEC PA_PNT
lbcr.20210209.024236.fits[2] Q0636+680 20 0 06:42:07.812 +67:58:35.61 90.46077
lbcr.20210209.024701.fits[2] Q0636+680 20 20 06:42:07.813 +67:58:55.60 90.19545
lbcr.20210209.025129.fits[2] Q0636+680 0 20 06:42:04.252 +67:58:55.64 90.21614

lbcr.20210209.030039.fits[2] Q0636+680 25 0 06:42:00.701 +67:58:55.60 90.20184 <-- PCSGUI did not show these offsets, but they are done
lbcr.20210209.030529.fits[2] Q0636+680 30 25 06:42:09.593 +67:59:00.60 90.33985
lbcr.20210209.031007.fits[2] Q0636+680 0 30 06:42:04.258 +67:59:05.61 90.26209

LBC_initial_5ptcustom_pa90.xml     
         <XDIT>20, 20, 0, -20, -20</XDIT>
         <YDIT>0, 20, 20, 20, 0</YDIT>
LBC_repeat_5ptcustom_pa90.xml
         <XDIT>25, 30, 0, -25, -30</XDIT>
         <YDIT>0, 25, 30, 25, 0</YDIT>

Active Optics test

John said this was done through the MODS1 collimation and that there is not a need to do MODS2. The MODS2 changes were trivial. We'll try nevertheless but will stop to reconfigure for PEPSI when necessary.

03:15 Reconfiguring to MODS+MODS.

03:28 David will check pointing and collimation on both MODS. Collimation on-axis. Works fine on both sides.

03:41 Reconfiguring to PEPSI-PFU. After pointing/collimation, Ilya is sending the preset at 09:04. Logging PEPSI observation in the google doc...

LBC hysteresis

Following these instructions, I created the catalog OSA0627. GS 11 is good and the ob file is created to take a series of 999 windowed (Y=2916-3915) images around it at V(LBCB) and r(LBCR).

10:19 Configured for the LBCs

10:20 David clicked collimate - we are on the collimation model

10:26 Slewing to collimation field (which has OSA627-11 on-axis).
  • Pointing offsets were non-zero and the M1s went into the limits. David removed these but they returned.
  • David restarted PCS, but MODE 3 pointing offsets are still there: 1.56, 1.904, 0, 40.9, -33.012, 0
  • Slewed again and the pointing offsets are non-zero, but smaller (tenths of a mm, rather than 1.5-2)
    • -0.461, -0.172, 0, -3.702, 9.908, 0 on SX and
    • -0.461, -0.143, 0.0, -3.075, 9.908, 0 on DX
10:39 dohybrid
  • LBCB chip 2 noise
10:48 Copointing images: LBCB, LBCR: 105007, 105000

10:52 Andrew set zero
  • 2 images with not telescope, not guiding
  • 2 images with not telescope, not guidings, not mirror.
  • start 999 images with not telescope, not guiding and mirror
11:00 there is definitely an x-shift and one star is on the edge on LBCB but not on the edge on LBCR
  • copoint using 105956, 105946 for LBCB, LBCR. The dCE, dCA = ~5". The difference is 9".
  • Start taking images.
11:07 Andrew re-referenced TMS

11:10 Slew to zenith (slew to hold 90 degl elevation) --- 11:11:40 at zenith

11:20:10 Slew to OSA0627-11 via newpoint

11:21:40 at target on both sides.
  • 112157, 112229
11:26:30 Slew to hold at 31.1 deg --- 11:28

11:33 Sent preset back to field -- arrived at field at 11:34:15

11:39 Slew to hold at zenith --- arrived at 11:40:30

11:45:30 Slew back to field -- arrived at 11:47:

  • Way out of focus on the blue side, initially.
11:52 Slew back to 31.1 deg --- arrived at 11:53

11:58 Slew back to field -- arrived at 11:59:06

12:04 Slewing back to zenith -- arrived at 12:05:34

12:10 Back to field -- arrived at 12:12:11

12:17 Slewing to 31.1 deg -- arrived at 12:18:30

12:23 Back to field --- arrived at 12:24:39

12:29:30 Continuing exposure while Andrew calculates adjustments needed to refocus/recollimate based on TMS. Continue taking images meanwhile.

And the answers are... 0.0894, -0.0544, -0.1251, -10.645", 1.363"

for z, -0.38 was already there, the value I entered was -0.5051

12:38:11 updated the global offsets

There was a sign error, so we are putting the global offsets back to what they were originally.

Now reverse the signs on the numbers above. So we now enter: -0.0894, 0.0544, -0.2549 10.6 -1.36

12:44 updated

12:46 Now entered Y global to 0.4 and updated.

Add 0.045 to X global --> -0.0444 updated at 12:48

To Y add 0.095, so Y global = 0.0495 and updated 12:49

Images 124931 for blue and 124927 for red are images before FPIA

Ran dofpia and it took 3 iterations to converge. The corrections will be noted here: (ended at 125509)
  Z4        Z5    Z6     Z7     Z8    Z11      Z22
 -1050      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  3796    -25   -191    265    446    -27      7    y  1.13  1.52                 /lbt/data/new/lbcb.20210209.125206.fits  V-BESSEL
   274   -171    -56    -85     27    202     48    y  1.66  0.20                 /lbt/data/new/lbcb.20210209.125343.fits  V-BESSEL
  -365    185    -54   -202   -236    147     70    y  1.69  0.24                 /lbt/data/new/lbcb.20210209.125515.fits  V-BESSEL
  1208      0      0      0      0   -690      0    y  0.00  0.00                     dofpia_sph3_adjustment             

 -875      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
 -3897    126     -3    128    -93    181     65    y  0.88  1.54                 /lbt/data/new/lbcr.20210209.125202.fits  r-SLOAN 
  -555     75    -21   -157     -1    140     44    y  0.95  0.26                 /lbt/data/new/lbcr.20210209.125338.fits  r-SLOAN 
    15    104   -143    308    -33     83     36    y  1.24  0.18                 /lbt/data/new/lbcr.20210209.125509.fits  r-SLOAN 
   888      0      0      0      0   -507      0    y  0.00  0.00                     dofpia_sph3_adjustment             

13:07 Slew to high elevation field (elevation 78:57), OSA0634 gs=0. Andrew notes some Y/RX coma.

13:10 Started dofpia (the global offsets that I had entered are still in place). The first pupils look good - FPIA converaged in 2 iterations.

13:18 Taking a few images (loading the "current" position OB from the previous field and playing it).

robs (or my x2go onto robs) is crashing - stalled and then came back to life. Same thing at 13:34. I killed all ds9s.

Try g+r flats from obs3, but it is too bright.

13:43 David is closing up. I restored teh Z global offsets on DX so all are zero except Z_go = -0.38mm.

13:44 robs came back to life.

    Create the catalog
        ssh telescope@obs1
        pngui
            Click CREATE tab
            Enter coordinates where you want to slew, then
            Click "Create Catalog Entry" and "SAVE". This will search for stars within "Err" of the input coordinates and create a catalog entry "OSA###.catalog" under /lbt/data/catalogs/lbto/Scratch/OSA/catalogs 
    To Slew to a star in the new catalog:
        open 2 IRAF sessions and set up LBTtools:
            camera = LBC (or RedLBC or BlueLBC)
            Telside = "Left" in one and "Right" in the other IRAF session 
        load LBTtools and Observe
        epar newpoint to enter the catalog name, TelMode and gstarID
        Issue the command, newpoint BC=1, on both sides. 

-- BarryRothberg - 07 Feb 2021
Topic revision: r6 - 09 Feb 2021, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback